OpenStudy (anonymous):

A 25 N object requires a 5.0 N to start moving over a horizontal surface. What is the coefficient of static friction?

OpenStudy (anonymous):

f = Us * N f/ N = Us

OpenStudy (anonymous):

This is what I got, is it correct? F = m*ks 25*0.2 = 5 ks=0.2